OpenStudy (anonymous):

How do you find the vertex and directrix in the equation 4/(1-sinθ)

OpenStudy (amistre64):

you seem to have a polar setup, recall that y = r sin(theta)

OpenStudy (anonymous):

I found that the directrix is y=-4 because of the equation

OpenStudy (amistre64):

im not familiar with the properties like that :) r = 4/(1-sin) r(1-sin) = 4 r-r sin = 4 r - y = 4 sqrt(x^2+y^2) - y = 4 sqrt(x^2+y^2) = 4 + y x^2 + y^2 = 16+8y + y^2 x^2 = 16+8y x^2 - 16 = 8y (x^2 - 16)8 = y

OpenStudy (amistre64):

missed the / when i divide by 8 :/

OpenStudy (amistre64):

so it looks to me like an upward parabola, symmetric about the y axis, but shifted down by 16

OpenStudy (anonymous):

do you think that I could find the vertex by plotting points?

OpenStudy (anonymous):

It has to be polar

OpenStudy (amistre64):

you could, but as is it looks to me from my manipulations that when x=0 y=-2

OpenStudy (amistre64):

so when r = sqrt(0^2+2^2) = 2 4/(1-sin) = 2 2 = 1-sin sin = -1 when theta = -pi/2

OpenStudy (anonymous):

ok thank you :)

OpenStudy (amistre64):

youre welcome ... with any luck those random synapse firings make a little sense :)